In the following method:

Reaction:

In separate flasks, prepare three solutions:

i) copper (II) acetate (1.7 g) in water (50 ml) (may require some warming to assist dissolving)

ii) salicaldehyde (2.7 ml) in methanol (50 ml) along with methylamine (5.2 ml, 33% in ethanol) and methanol (50 ml) and stand in a covered conical flask (500 ml) at room temperature for 5 minutes.

iii) dissolve sodium acetate (5.0 g) in water (50 ml).

Place a magnetic stirrer bean into the 500 ml conical flask and place the flask on a magnetic stirrer. Transfer the copper acetate and sodium acetate solutions into two separate separating funnels and with efficient stirring, add the copper acetate and sodium acetate solutions dropwise TO the salicaldehyde solution over 10 minutes at the same time. Heat the resulting solution to flocculation (approx. 50˚ C) for 10 minutes with occasional stirring.

Workup: Cool the solution to room temperature by immersing the flask in an ice bath and remove the precipitated material by filtration. Wash the material with water (3 x 15 ml) and cold ethanol (5 ml).

Purification: Recrystalise the dark green product from methanol (~140 ml) and allow to air dry.

Is the reaction between salicaldehyde and methyl amine to form a Schiff's base a 1:1 ratio?

And what would the theoretical yield be?
Answer #1

Salicylaldehyde = 2.7 ml

Density = 1.146 g/ml

Methyl amine in 33% methanol = 5.2 ml

Density of methylamine in 33% methanol = 0.756 g/ml

   CH3NH2 + C6H4(OH)CHO ----------> CH3 - N = HC-(OH)-C6H4

Methyl amine Salicylaldehyde Schiff base

From the above reaction we can observe that molar ratio of methylamine and salicylaldehyde = 1:1

Mass of Salicylaldehyde = Density x volume = 1.146 g/ml x 2.7 ml = 3.0942 g

Molar mass of salicylaldehyde = 122.12 g/mol

Mass of methyl amine = Density x volume = 0.756 g/ml x 5.2 ml = 3.9312 g

Molar mass of methyl amine = 31.05 g/mol

No. of moles of salicylaldehyde = mass/molar mass = 3.0942 g/122.12 g/mol = 0.025 moles

No. of moles of methyl amine = 3.9312 g/31.05 g/mol = 0.127 moles

From the above two reactants we can find that salicylaldehyde is limiting reactant

Molar ratio of salicylaldehyde and Schiff's base = 1:1

Hence theoretical moles of Schiff's base = 0.025 moles

molar mass of Schiff's base = 135.08 g/mol (8x12.01(C)+ 9x1(H) + 16x1(O)+14x1(N))

Theoretical yield of Schiff's base = molar mass x theoretical moles = 135.08 g/mol x 0.025 mol = 3.377 g